OpenStudy (peterg2001):

help will medal All of the following equations have the same solution except _____. 3m + 2 = -10 2 - 3m = -10 -10 = -3m + 2 -3m + 2 = -10

OpenStudy (sagewilson):

I think it's A but I'm not so sure

OpenStudy (lcnettech):

3m + 2 = -10

OpenStudy (lcnettech):

so a would be your answer

OpenStudy (jj123.3):

its A

OpenStudy (anonymous):

a they all equal 4 except for a it equals -4

OpenStudy (anonymous):

yes I am sure

OpenStudy (sagewilson):

so A

OpenStudy (anonymous):

http://www.mathpapa.com/

OpenStudy (anonymous):

3m + 2 = -10 = -4 2 -3m = -10 = 4 -10 = -3m + 2 = 4 -3m + 2 = -10 = 4 Like everyone said it's A